# Finding Safety Violations of AI-Enabled Control Systems through the Lens of Synthesized Proxy Programs

### v1

@@ 1. Introduction
Control systems, which utilize software controllers to maintain desired behaviors of particular mechanical systems, have been broadly deployed into diverse real-world applications. These systems have become an essential part of everyday life and industrial production (Xia et al., 2015). With the rapid advances of Artificial Intelligence (AI) in recent decades, there has been a huge demand for enhancing or even replacing traditional mathematical controllers (e.g., Proportional-Integral-Derivative (Johnson and Moradi, 2005)) with AI controllers (e.g., Deep Neural Networks) to achieve more optimized performance. We refer to control systems employing AI controllers as AI-enabled control systems. With their immense potential, AI-enabled control systems have inevitably been or will soon be adopted in various safety-critical domains, such as aerospace (Izzo et al., 2019; Chai et al., 2021), autonomous driving (Kiran et al., 2021; Muhammad et al., 2020), healthcare monitoring (Kishor and Chakraborty, 2022; Kamruzzaman, 2021), and even nuclear fusion (Degrave et al., 2022). Therefore, ensuring the reliability and safety of AI-enabled control systems in such contexts is of paramount importance.
To date, falsification has been shown to be an effective approach for testing complex control systems (Song et al., 2022; Corso et al., 2021; Menghi et al., 2020; Yaghoubi and Fainekos, 2017; Zhang et al., 2022, 2021a). Concretely, given a control system with a safety specification, usually written in Signal Temporal Logic (STL) (Maler and Nickovic, 2004), falsification aims to find an input signal that causes the control system to violate the specification. Falsification often uses an optimization algorithm to search for input signals with the goal of minimizing the quantitative robustness semantics of temporal logics (Donzé and Maler, 2010). These semantics quantify the degree to which an input signal deviates from violating the safety specification as a real number. A smaller number means that the corresponding input signal is more likely to violate a given safety specification, and a negative number suggests that the safety specification is violated. Falsification iteratively generates candidate input signals and executes them on the control system until an input with negative robustness semantics is found, revealing a violation of the safety specification. Various falsification methods have been proposed and adopted in practice (Eddeland et al., 2020; Tuncali et al., 2018; Kapinski et al., 2016), showing that falsification is effective in revealing specification violations in complex control systems.
@@
Second, falsification of AI-enabled control systems fails to comprehensively cover conjunctive specifications. AI-enabled control systems emerge to handle highly complex control tasks, often with multiple safety specifications. For instance, a self-driving control system in (Zhu et al., 2019) is required to maintain a car’s heading angle within $90^{\circ}$ ($\varphi_{1}$) while keeping the distance between the car and the centerline of the road below $2.0$ meters ($\varphi_{2}$). A common practice for falsifying such multiple specifications simultaneously is to combine them as a conjunctive specification ($\varphi_{1}\wedge\varphi_{2}$(Lidén Eddeland et al., 2022). However, directly falsifying such a conjunctive specification often produces less diverse safety violations. If one sub-specification is easily violated, the falsification process tends to repeatedly exploit that sub-specification to report violations and overlook others. As Sun et al. (Sun et al., 2022) discussed, it is desirable to find diverse violations that can cover as many sub-specifications as possible. Nevertheless, the existing state-of-the-art falsification tools like PSY-TaLiRo (Thibeault et al., 2021) are not inherently designed to meet this challenge.
@@
Additionally, we consider that recent studies have highlighted the importance of the exploration-exploitation trade-off in falsifying conjunctive specifications (Mathesen et al., 2021; Ramezani et al., 2022). Specifically, falsification can either exploit a sub-specification that is easily violated to ensure successful falsification, or explore a sub-specification that has not yet been falsified to boost the coverage of all sub-specifications within the conjunctive specification. The low coverage of sub-specifications is often a result of inadequate exploration and over-exploitation. Thus, we propose a strategy that balances the exploration and exploitation of sub-specifications, allowing falsification to more comprehensively cover the conjunctive specification.
@@ 2.1. AI-enabled Control System
Figure 1 shows a typical architecture of AI-enabled control systems consisting of a plant, a controller, sensors, and actuators. The plant $\mathcal{M}$ is a physical environment whose next state $y^{\prime}$ is determined by the current state $y$ and the control signal $u$. The control signal $u$ is issued by an AI controller and applied to the plant by actuators. Integrating AI controllers allows control systems to adapt to more complex situations and often achieve superior results compared to traditional methods (Hunt et al., 1992; Lillicrap et al., 2016). However, this integration also introduces new risks, such as the potential for system failures that could violate safety requirements. These risks underscore the need for our work in analyzing the safety and reliability of AI-enabled control systems.
The AI controller is typically implemented as a deep neural network, and Deep Reinforcement Learning (DRL) is currently considered the state-of-the-art method for training it (Song et al., 2022; Duan et al., 2016; Lillicrap et al., 2016). Our study focuses on control systems with DRL-based AI controllers. DRL is a subfield of machine learning that focuses on learning optimal decision strategies through interactions with the environment (Wang et al., 2022). In DRL, an agent learns the best policy by observing states and selecting the best output to maximize the cumulative reward. Concretely, considering a DRL-based AI controller as an agent, it receives the current state of the plant and outputs a control signal to transition the system to a new state. A policy, a deep neural network taking the system’s state as input, determines the control signal that the agent outputs. The reward for each output under a state depends on the feedback from the plant. Through repeated interactions with the plant, the agent is trained to maximize the reward it receives, learning and adapting its policy to make better decisions over time. After training, the agent with its best policy is deployed as an AI controller in the control system. DRL-based AI controllers have found applications in many critical areas such as autonomous driving (Kiran et al., 2021; Wu et al., 2021), robotics control (Ju et al., 2022; Hester et al., 2012), etc.
@@ 2.4. $\epsilon$-greedy Strategy
When iteratively selecting among multiple options, a key challenge is to balance exploration (i.e., trying new options that might yield better results) and exploitation (i.e., choosing the best option based on current knowledge). In this context, the $\epsilon$-greedy strategy (Sutton and Barto, 2018) offers a straightforward yet effective solution. The $\epsilon$ in its name represents a probability that governs the trade-off between exploration and exploitation. Formally, consider a set of $n$ options $O=\{o_{1},o_{2},\dots,o_{n}\}$, where the performance of each option is evaluated through reward feedback, we can obtain the knowledge of the expected reward =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ of each option by averaging the rewards received from multiple attempts so far. Then, as shown in Algorithm 2, with a probability of $1-\epsilon$, the strategy exploits the current knowledge by selecting the option with the highest expected reward (line 2). Alternatively, with probability $\epsilon$, it explores by selecting an option at random (line 2), enabling the discovery of potentially better alternatives.
@@
As shown by Zhang et al. (Zhang et al., 2021c), certain sub-specifications within a conjunctive specification are significantly easier to falsify than others as they are more sensitive to input signals. This sensitivity causes falsification algorithms to disproportionately focus on these easier sub-specifications, reducing the likelihood of falsifying more difficult ones and potentially resulting in an incomplete safety analysis. Our approach treats each sub-specification as an independent option and incorporates the $\epsilon$-greedy strategy to encourage the falsification of sub-specifications that have not yet been falsified, thereby ensuring a more comprehensive safety analysis. A detailed explanation of how the $\epsilon$-greedy strategy works in our approach is provided in Section 4.2.
@@ 3. Motivating Example
Key Challenges.   We use the state-of-the-art falsification tool, PSY-TaLiRo (Thibeault et al., 2021), which employs the Simulated Annealing-based falsification algorithm outlined in Algorithm 1, to find safety violations of the self-driving system. We set the maximum number of falsification trials to 50 and feed the above-mentioned safety specification into PSY-TaLiRo. The results of running time and sub-specification violation coverage are presented in Table 1. Following Sun et al. (Sun et al., 2022), the sub-specification violation coverage is defined as follows:
@@ 4.1.2. Evolution Strategy
Once we have constructed the sketch, we then employ an Evolution Strategy (ES)-based synthesis algorithm to synthesize the coefficients in the sketch. ES is a well-known optimization algorithm that has been successfully applied to solve program synthesis (Sobania et al., 2022) and controller synthesis problems (Salimans et al., 2017). The ES-based synthesis algorithm is presented in Algorithm 3. The algorithm first randomly initializes the coefficients $\theta$ in the sketch (line 3), then performs the following steps iteratively:
@@ 5.1.1. Benchmark
We utilized the well-trained AI controllers for the 8 control systems from the official implementation of Zhu et al. (Zhu et al., 2019), where these controllers were trained by Deep Deterministic Policy Gradient (Lillicrap et al., 2016), a popular DRL algorithm for continuous control tasks. Table 2 shows the numbers of input and output variables of the AI controllers as well as their neural network sizes. The AI controllers have a wide range of sizes, featuring 2 or 3 layers with neuron counts varying from 440 (240+200) to 1200 (500+300+400); such sizes are sufficiently large for AI controllers (e.g., those controllers used in (Zhang et al., 2022; Tran et al., 2020) have no more than 400 neurons).
@@ 5.1.2. Baseline
In our evaluation, we did not include several other open-source black-box falsification tools, such as Breach (Donzé, 2010), ARIsTEO (Menghi et al., 2020), FalStar (Ernst et al., 2021b), ForeSee (Zhang et al., 2021b), zlscheck (Ismailbennani, [n.d.]), NNV (Tran et al., 2020), and Mosaic (Xie et al., 2023). These tools are specialized for MATLAB/Simulink models and written in different programming languages such as MATLAB, Java, Scala, or OCaml. The features of different programming languages can greatly impact the experimental results. Also, integrating these tools into our evaluation could pose significant compatibility challenges with the Python-developed control systems and toolchains used in our experiments. We believe that comparing Synthify to PSY-TaLiRo, the only existing falsification toolbox natively in Python, is fair and appropriate since PSY-TaLiRo is a state-of-the-art and well-established falsification tool that has been well studied. The 2022 ARCH-COMP prize awarded to PSY-TaLiRo also demonstrates its advancement and maturity in falsification (Ernst et al., 2022).
@@ 5.2. Implementation and Hyperparameters
Synthify is built on top of the implementation of Simulated Annealing (SA)-based falsification algorithm and other infrastructure components from PSY-TaLiRo (version 1.0.0b3). To compute the quantitative robustness semantics, we integrate an off-the-shelf tool called RTAMT (Ničković and Yamaguchi, 2020) (version 0.3.0) into our framework. We set the maximum number of iterations for the SA-based algorithm to 100, a recommended value by recent literature (Menghi et al., 2020; Yamagata et al., 2020). By reusing all other default hyperparameters of PSY-TaLiRo in the SA-based algorithm, we ensure a fair comparison while also evaluating the unique design of Synthify. For the Evolution Strategy (ES)-based synthesis algorithm, we use a population size of 50, and both the maximum number of iterations and time steps are set to 100. The noise standard deviation and the learning rate are set to 0.1 and 0.05, respectively. The threshold of the fitness is set to 0.1. The refinement follows the same hyperparameters as the ES-based synthesis algorithm, except that the learning rate is set to 1e-3 to avoid overfitting. In our $\epsilon$-greedy sampling strategy, we set the probability threshold $\epsilon$ to 0.9. All the experiments are conducted using an Ubuntu 18.04 server equipped with an Intel Xeon E5-2698 CPU and 504 GB RAM.
@@ 5.3. Experimental Setup and Metrics
  • Setting 2: Following the common practice of search-based testing (Kang et al., 2022; Pang et al., 2022; Zolfagharian et al., 2023), we specify the budget as a fixed duration of time, which is set to 10 minutes, in line with prior studies (Kang et al., 2022; Soltani et al., 2018).
@@
  • Setting 1: The effectiveness of each tool is evaluated by the falsification success rate, which is the ratio of the number of trials that uncover violations to 50 trials. This metric is commonly adopted in falsification studies (Ernst et al., 2022, 2021a; Zhang et al., 2022). The efficiency of each tool is measured by the falsification time, which is the time required to complete 50 falsification trials.
  • Setting 2: The effectiveness of each tool is evaluated using the metric of the number of violations found by each tool within the given time budget of 10 minutes, following prior search-based testing studies (Sun et al., 2022; Kang et al., 2022). The efficiency of each tool is measured by the time required to reveal one safety violation.
@@ 6.1. RQ1: Effectiveness of Synthify
We also use the Vargha-Delaney statistic $\hat{A}_{12}$ (Vargha and Delaney, 2000) to measure the effect size. This non-parametric measure gives the probability that a randomly chosen value from one group is higher or lower than one from another group. It is commonly used to evaluate search-based testing methods due to their randomness (Palomba et al., 2016; Meng et al., 2022; Olsthoorn et al., 2020). According to Vargha and Delaney (Vargha and Delaney, 2000), $\hat{A}_{12}$ values greater than 0.71 (or less than 0.29) indicate a “large” effect size. Values between 0.64 and 0.71 (or between 0.29 and 0.36) correspond to a “medium” effect size, while values between 0.36 and 0.64 suggest a “negligible” effect size. As shown in Table 3, Synthify significantly outperforms PSY-TaLiRo with a large effect size on 6 out of 8 control systems. Only on CartPole and Quadcopter, we observe negligible and medium effect sizes. But overall, the results show that Synthify is more effective than PSY-TaLiRo in finding violations of safety specifications.

### v2

@@ 1. Introduction
Control systems, which utilize software controllers to maintain desired behaviors of particular mechanical systems, have been broadly deployed into diverse real-world applications. These systems have become an essential part of everyday life and industrial production (Xia et al., 2015). With the rapid advances of Artificial Intelligence (AI) in recent decades, there has been a huge demand for enhancing or even replacing traditional mathematical controllers (e.g., Proportional-Integral-Derivative (Johnson and Moradi, 2005)) with AI controllers (e.g., Deep Neural Networks) to achieve more optimized performance. We refer to control systems employing AI controllers as AI-enabled control systems. With their immense potential, AI-enabled control systems have inevitably been or will soon be adopted in various safety-critical domains, such as aerospace (Izzo et al., 2019; Chai et al., 2021), autonomous driving (Kiran et al., 2022; Muhammad et al., 2021), healthcare monitoring (Kishor and Chakraborty, 2022; Kamruzzaman, 2021), and even nuclear fusion (Degrave et al., 2022). Therefore, ensuring the reliability and safety of AI-enabled control systems in such contexts is of paramount importance.
To date, falsification has been shown to be an effective approach for testing complex control systems (Song et al., 2022; Corso et al., 2022; Menghi et al., 2020; Yaghoubi and Fainekos, 2017; Zhang et al., 2023a, 2022). Concretely, given a control system with a safety specification, usually written in Signal Temporal Logic (STL) (Maler and Nickovic, 2004), falsification aims to find an input signal that causes the control system to violate the specification. Falsification often uses an optimization algorithm to search for input signals with the goal of minimizing the quantitative robustness semantics of temporal logics (Donzé and Maler, 2010). These semantics quantify the degree to which an input signal deviates from violating the safety specification as a real number. A smaller number means that the corresponding input signal is more likely to violate a given safety specification, and a negative number suggests that the safety specification is violated. Falsification iteratively generates candidate input signals and executes them on the control system until an input with negative robustness semantics is found, revealing a violation of the safety specification. Various falsification methods have been proposed and adopted in practice (Eddeland et al., 2020; Tuncali et al., 2018; Kapinski et al., 2016), showing that falsification is effective in revealing specification violations in complex control systems.
@@
Second, falsification of AI-enabled control systems fails to comprehensively cover conjunctive specifications. AI-enabled control systems emerge to handle highly complex control tasks, often with multiple safety specifications. For instance, a self-driving control system in (Zhu et al., 2019) is required to maintain a car’s heading angle within $90^{\circ}$ ($\varphi_{1}$) while keeping the distance between the car and the centerline of the road below $2.0$ meters ($\varphi_{2}$). A common practice for falsifying such multiple specifications simultaneously is to combine them as a conjunctive specification ($\varphi_{1}\wedge\varphi_{2}$(Lidén Eddeland et al., 2022). However, directly falsifying such a conjunctive specification often produces less diverse safety violations. If one sub-specification is easily violated, the falsification process tends to repeatedly exploit that sub-specification to report violations and overlook others. As Sun et al. (Sun et al., 2023) discussed, it is desirable to find diverse violations that can cover as many sub-specifications as possible. Nevertheless, the existing state-of-the-art falsification tools like PSY-TaLiRo (Thibeault et al., 2021) are not inherently designed to meet this challenge.
@@
Additionally, we consider that recent studies have highlighted the importance of the exploration-exploitation trade-off in falsifying conjunctive specifications (Mathesen et al., 2021; Ramezani et al., 2023). Specifically, falsification can either exploit a sub-specification that is easily violated to ensure successful falsification, or explore a sub-specification that has not yet been falsified to boost the coverage of all sub-specifications within the conjunctive specification. The low coverage of sub-specifications is often a result of inadequate exploration and over-exploitation. Thus, we propose a strategy that balances the exploration and exploitation of sub-specifications, allowing falsification to more comprehensively cover the conjunctive specification.
@@ 2.1. AI-enabled Control System
Figure 1 shows a typical architecture of AI-enabled control systems consisting of a plant, a controller, sensors, and actuators. The plant $\mathcal{M}$ is a physical environment whose next state $y^{\prime}$ is determined by the current state $y$ and the control signal $u$. The control signal $u$ is issued by an AI controller and applied to the plant by actuators. Integrating AI controllers allows control systems to adapt to more complex situations and often achieve superior results compared to traditional methods (Hunt et al., 1992; Lillicrap et al., 2019). However, this integration also introduces new risks, such as the potential for system failures that could violate safety requirements. These risks underscore the need for our work in analyzing the safety and reliability of AI-enabled control systems.
The AI controller is typically implemented as a deep neural network, and Deep Reinforcement Learning (DRL) is currently considered the state-of-the-art method for training it (Song et al., 2022; Duan et al., 2016; Lillicrap et al., 2019). Our study focuses on control systems with DRL-based AI controllers. DRL is a subfield of machine learning that focuses on learning optimal decision strategies through interactions with the environment (Wang et al., 2024a). In DRL, an agent learns the best policy by observing states and selecting the best output to maximize the cumulative reward. Concretely, considering a DRL-based AI controller as an agent, it receives the current state of the plant and outputs a control signal to transition the system to a new state. A policy, a deep neural network taking the system’s state as input, determines the control signal that the agent outputs. The reward for each output under a state depends on the feedback from the plant. Through repeated interactions with the plant, the agent is trained to maximize the reward it receives, learning and adapting its policy to make better decisions over time. After training, the agent with its best policy is deployed as an AI controller in the control system. DRL-based AI controllers have found applications in many critical areas such as autonomous driving (Kiran et al., 2022; Wu et al., 2023), robotics control (Ju et al., 2022; Hester et al., 2012), etc.
@@ 2.4. $\epsilon$-greedy Strategy
When iteratively selecting among multiple options, a key challenge is to balance exploration (i.e., trying new options that might yield better results) and exploitation (i.e., choosing the best option based on current knowledge). In this context, the $\epsilon$-greedy strategy (Sutton and Barto, 1998) offers a straightforward yet effective solution. The $\epsilon$ in its name represents a probability that governs the trade-off between exploration and exploitation. Formally, consider a set of $n$ options $O=\{o_{1},o_{2},\dots,o_{n}\}$, where the performance of each option is evaluated through reward feedback, we can obtain the knowledge of the expected reward =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ of each option by averaging the rewards received from multiple attempts so far. Then, as shown in Algorithm 2, with a probability of $1-\epsilon$, the strategy exploits the current knowledge by selecting the option with the highest expected reward (line 2). Alternatively, with probability $\epsilon$, it explores by selecting an option at random (line 2), enabling the discovery of potentially better alternatives.
@@
As shown by Zhang et al. (Zhang et al., 2021b), certain sub-specifications within a conjunctive specification are significantly easier to falsify than others as they are more sensitive to input signals. This sensitivity causes falsification algorithms to disproportionately focus on these easier sub-specifications, reducing the likelihood of falsifying more difficult ones and potentially resulting in an incomplete safety analysis. Our approach treats each sub-specification as an independent option and incorporates the $\epsilon$-greedy strategy to encourage the falsification of sub-specifications that have not yet been falsified, thereby ensuring a more comprehensive safety analysis. A detailed explanation of how the $\epsilon$-greedy strategy works in our approach is provided in Section 4.2.
@@ 3. Motivating Example
Key Challenges.   We use the state-of-the-art falsification tool, PSY-TaLiRo (Thibeault et al., 2021), which employs the Simulated Annealing-based falsification algorithm outlined in Algorithm 1, to find safety violations of the self-driving system. We set the maximum number of falsification trials to 50 and feed the above-mentioned safety specification into PSY-TaLiRo. The results of running time and sub-specification violation coverage are presented in Table 1. Following Sun et al. (Sun et al., 2023), the sub-specification violation coverage is defined as follows:
@@ 4.1.2. Evolution Strategy
Once we have constructed the sketch, we then employ an Evolution Strategy (ES)-based synthesis algorithm to synthesize the coefficients in the sketch. ES is a well-known optimization algorithm that has been successfully applied to solve program synthesis (Sobania et al., 2023) and controller synthesis problems (Salimans et al., 2017). The ES-based synthesis algorithm is presented in Algorithm 3. The algorithm first randomly initializes the coefficients $\theta$ in the sketch (line 3), then performs the following steps iteratively:
@@ 5.1.1. Benchmark
We utilized the well-trained AI controllers for the 8 control systems from the official implementation of Zhu et al. (Zhu et al., 2019), where these controllers were trained by Deep Deterministic Policy Gradient (Lillicrap et al., 2019), a popular DRL algorithm for continuous control tasks. Table 2 shows the numbers of input and output variables of the AI controllers as well as their neural network sizes. The AI controllers have a wide range of sizes, featuring 2 or 3 layers with neuron counts varying from 440 (240+200) to 1200 (500+300+400); such sizes are sufficiently large for AI controllers (e.g., those controllers used in (Zhang et al., 2023a; Tran et al., 2020) have no more than 400 neurons).
@@ 5.1.2. Baseline
In our evaluation, we did not include several other open-source black-box falsification tools, such as Breach (Donzé, 2010), ARIsTEO (Menghi et al., 2020), FalStar (Ernst et al., 2021b), ForeSee (Zhang et al., 2021a), zlscheck (Ismailbennani, [n.d.]), NNV (Tran et al., 2020), and Mosaic (Xie et al., 2023). These tools are specialized for MATLAB/Simulink models and written in different programming languages such as MATLAB, Java, Scala, or OCaml. The features of different programming languages can greatly impact the experimental results. Also, integrating these tools into our evaluation could pose significant compatibility challenges with the Python-developed control systems and toolchains used in our experiments. We believe that comparing Synthify to PSY-TaLiRo, the only existing falsification toolbox natively in Python, is fair and appropriate since PSY-TaLiRo is a state-of-the-art and well-established falsification tool that has been well studied. The 2022 ARCH-COMP prize awarded to PSY-TaLiRo also demonstrates its advancement and maturity in falsification (Ernst et al., 2022).
@@ 5.2. Implementation and Hyperparameters
Synthify is built on top of the implementation of Simulated Annealing (SA)-based falsification algorithm and other infrastructure components from PSY-TaLiRo (version 1.0.0b3). To compute the quantitative robustness semantics, we integrate an off-the-shelf tool called RTAMT (Ničković and Yamaguchi, 2020) (version 0.3.0) into our framework. We set the maximum number of iterations for the SA-based algorithm to 100, a recommended value by recent literature (Menghi et al., 2020; Yamagata et al., 2021). By reusing all other default hyperparameters of PSY-TaLiRo in the SA-based algorithm, we ensure a fair comparison while also evaluating the unique design of Synthify. For the Evolution Strategy (ES)-based synthesis algorithm, we use a population size of 50, and both the maximum number of iterations and time steps are set to 100. The noise standard deviation and the learning rate are set to 0.1 and 0.05, respectively. The threshold of the fitness is set to 0.1. The refinement follows the same hyperparameters as the ES-based synthesis algorithm, except that the learning rate is set to 1e-3 to avoid overfitting. In our $\epsilon$-greedy sampling strategy, we set the probability threshold $\epsilon$ to 0.9. All the experiments are conducted using an Ubuntu 18.04 server equipped with an Intel Xeon E5-2698 CPU and 504 GB RAM.
@@ 5.3. Experimental Setup and Metrics
  • Setting 2: Following the common practice of search-based testing (Kang et al., 2022; Pang et al., 2022; Zolfagharian et al., 2023), we specify the budget as a fixed duration of time, which is set to 10 minutes, in line with prior studies (Kang et al., 2022; Soltani et al., 2020).
@@
  • Setting 1: The effectiveness of each tool is evaluated by the falsification success rate, which is the ratio of the number of trials that uncover violations to 50 trials. This metric is commonly adopted in falsification studies (Ernst et al., 2022, 2021a; Zhang et al., 2023a). The efficiency of each tool is measured by the falsification time, which is the time required to complete 50 falsification trials.
  • Setting 2: The effectiveness of each tool is evaluated using the metric of the number of violations found by each tool within the given time budget of 10 minutes, following prior search-based testing studies (Sun et al., 2023; Kang et al., 2022). The efficiency of each tool is measured by the time required to reveal one safety violation.
@@ 6.1. RQ1: Effectiveness of Synthify
We also use the Vargha-Delaney statistic $\hat{A}_{12}$ (Vargha and Delaney, 2000) to measure the effect size. This non-parametric measure gives the probability that a randomly chosen value from one group is higher or lower than one from another group. It is commonly used to evaluate search-based testing methods due to their randomness (Palomba et al., 2016; Meng et al., 2022; Olsthoorn et al., 2021). According to Vargha and Delaney (Vargha and Delaney, 2000), $\hat{A}_{12}$ values greater than 0.71 (or less than 0.29) indicate a “large” effect size. Values between 0.64 and 0.71 (or between 0.29 and 0.36) correspond to a “medium” effect size, while values between 0.36 and 0.64 suggest a “negligible” effect size. As shown in Table 3, Synthify significantly outperforms PSY-TaLiRo with a large effect size on 6 out of 8 control systems. Only on CartPole and Quadcopter, we observe negligible and medium effect sizes. But overall, the results show that Synthify is more effective than PSY-TaLiRo in finding violations of safety specifications.